# Interpreting Graph Neural Networks with In-Distributed Proxies

### v1

@@ Abstract
Graph Neural Networks (GNNs) have become a building block in graph data processing, with wide applications in critical domains. The growing needs to deploy GNNs in high-stakes applications necessitate explainability for users in the decision-making processes. A popular paradigm for the explainability of GNNs is to identify explainable subgraphs by comparing their labels with the ones of original graphs. This task is challenging due to the substantial distributional shift from the original graphs in the training set to the set of explainable subgraphs, which prevents accurate prediction of labels with the subgraphs. To address it, in this paper, we propose a novel method that generates proxy graphs for explainable subgraphs that are in the distribution of training data. We introduce a parametric method that employs graph generators to produce proxy graphs. A new training objective based on information theory is designed to ensure that proxy graphs not only adhere to the distribution of training data but also preserve essential explanatory factors. Such generated proxy graphs can be reliably used for approximating the predictions of the true labels of explainable subgraphs. Empirical evaluations across various datasets demonstrate our method achieves more accurate explanations for GNNs.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a pivotal technology for handling graph-structured data, demonstrating remarkable performance in various applications including node classification and link prediction (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Scarselli et al., 2008). Their growing use in critical sectors such as healthcare and fraud detection has escalated the need for explainability in their decision-making processes (Zhang et al., 2022a; Wu et al., 2022; Li et al., 2022). To meet this demand, a variety of explanation methods have been recently developed to interpret the behavior of GNN models. These methods primarily concentrate on identifying a subgraph that significantly impacts the model’s prediction for a particular instance (Ying et al., 2019; Luo et al., 2020).
@@
Figure 1: Examples of the explanation process and out-of-distribution problem. (a) is the explanation process for a graph learning model. The original graph $G$ undergoes an explanation process, resulting in an explanation graph $G^{\prime}$ that highlights the most significant features and relationships; (b) shows the explanation graph $G^{\prime}$ is out of distribution where the GNN is well-trained.
Although it is intuitively correct, the underlying assumption of the aforementioned approach – GNN model can make accurate predictions on explanation subgraphs – may not always hold. Prior research indicates that explanation subgraphs can significantly deviate from the distribution of original graphs, leading to an Out-Of-Distribution (OOD) issue (Zhang et al., 2023b; Fang et al., 2023d, b). For instance, in the MUTAG dataset (Debnath et al., 1991), each graph represents a molecule, with nodes symbolizing atoms and edges indicating chemical bonds. The molecular graphs in this dataset usually contain hundreds of edges. In contrast, the $NO_{2}$ functional group, identified as a key subgraph influencing positive mutagenicity in a molecule, comprises merely 2 edges. This stark contrast in structural properties leads to a significant difference in the distributions of explanation subgraphs and original graphs. Since the model was trained with original graphs, the reliability of predictions on subgraphs is undermined due to the distribution shifting problem.
Several pioneering studies have attempted to address this distributional challenge (Fang et al., 2023d; Zhang et al., 2023b, a). For example, CGE regards the GNN model as a transparent, fully accessible system. It considers the GNN model as a teacher network and employs an additional “student” network to predict the labels of explanation subgraphs (Fang et al., 2023d). As another example, MixupExplainer (Zhang et al., 2023b) generates a mixed graph for the explanation by blending it with a non-explanatory subgraph from a different input graph. This method posits that the mixup graph aligns with the distribution of the original input graphs. However, this claim is predicated on a rather simplistic assumption that the explanation and non-explanatory subgraphs are independently drawn. However, in real-world applications, these methods often face practical limitations. The dependence on a “white box” model in CGE and the oversimplified assumptions in MixupExplainer are not universally applicable. Instead, GNN models are usually given as “black boxes”, and the graphs in real-life applications do not conform to strict independence constraints, highlighting the need for more versatile and realistic approaches to the OOD problem.
@@
  • We systematically analyze and address the challenge of the out-of-distribution issue in explainable GNNs, which is pivotal for enhancing the reliability and interpretability of GNNs in real-world applications.
@@
  • Through comprehensive experimentation and rigorous evaluations on various real-world datasets, we substantiate the effectiveness of our proposed approach, showcasing its practical utility and superiority in producing explanations.
@@ 2 Notations and Preliminary
2.1 Notations and Problem Formulation.
We denote a graph $G$ from an alphabet ${\mathcal{G}}$ by a triple $({\mathcal{V}},{\mathcal{E}};{\bm{X}})$, where ${\mathcal{V}}=\{v_{1},v_{2},...,v_{n}\}$ is the node set and ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the edge set. ${\bm{X}}\in{\mathbb{R}}^{n\times d}$ is the node feature matrix, where $d$ is the feature dimension and the $i$-th row is the feature vector associated with node $v_{i}$. The adjacency matrix of $G$ is denoted by ${\bm{A}}\in\{0,1\}^{n\times n}$, which is determined by the edge set $\mathcal{E}$ that ${A}_{ij}=1$ if $(v_{i},v_{j})\in\mathcal{E}$, ${A}_{ij}=0$, otherwise. In this paper, we focus on the graph classification task, as the node classification can be converted to computation graph classification problem (Ying et al., 2019; Luo et al., 2020). Specifically, for the graph classification task, each graph $G$ is associated with a label $Y\in\mathcal{C}$. The to-be-explained GNN model $f(\cdot)$ has been well-trained to classify $G$ into its class, i.e., $f:{\mathcal{G}}\mapsto\{1,2,\cdots,|\mathcal{C}|\}$.
Following existing works (Yuan et al., 2022; Luo et al., 2020; Ying et al., 2019), the explanation methods under consideration in this paper are model/task agnostic and treat GNN models as black boxes — i.e., the so-called post-hoc, instance-level explanation methods. Formally, our research problem is described as follows (Huang et al., 2024):
@@
Given a to-be-explained GNN model $f(\cdot)$ and a set of graphs ${\mathcal{G}}$, the goal of post-hoc instance-level explanation is to learn a parametric function that for an arbitrary graph $G\in{\mathcal{G}}$, it finds a compact subgraph $G^{*}\subseteq G$ that can “explain” the prediction $f(G)$. The parametric mapping $\Psi_{\psi}:{\mathcal{G}}\mapsto{\mathcal{G}}^{*}$ is called an explanation function, where ${\mathcal{G}}^{*}$ is the alphabet of $G^{*}$, and $\psi$ is the parameter of the explanation function.
2.2 Graph Information Bottleneck as the Objective Function.
The Information Bottleneck (IB) principle, foundational in learning dense representations, suggests that optimal representations should balance minimal and sufficient information for predictions (Tishby et al., 2000). This concept has been adapted for GNNs through the Graph Information Bottleneck (GIB) approach, consolidating various post-hoc GNN explanation methods like GNNExplainer (Ying et al., 2019) and PGExplainer (Luo et al., 2020). The GIB framework aim to find a subgraph $G^{*}$ to explain the GNN model’s prediction on a graph $G$ as (Zhang et al., 2023b):
@@
where $G^{\prime}$ is a candidate explanatory subgraph, $Y$ is the label, and $\alpha$ is a balance parameter. The first term is the mutual information between the original graph $G$ and the explanatory subgraph $G^{\prime}$ and the second term is negative mutual information of $Y$ and the explanation subgraph $G^{\prime}$. At a high level, the first term encourages detecting a small and dense subgraph for explanation, and the second term requires that the explanation subgraph is label preserving.
Due to the intractability of mutual information between $Y$ and $G^{\prime}$ in equation 1, existing works (Miao et al., 2022; Wu et al., 2020) derive a parameterized variational lower bound of $I(Y,G^{\prime})$:
@@
The OOD Problem in GIB. In existing research, the estimation of $P(Y|G^{\prime})$ is typically achieved by applying the to-be-explained model $f$, to the input graph $G^{\prime}$ (Ying et al., 2019; Miao et al., 2022). A critical assumption in existing approaches involves the GNN model’s ability to accurately predict candidate explanation subgraphs $G^{\prime}$. This assumption, however, often overlooks the OOD problem, where the distribution of explanation subgraphs significantly deviates from that of the original training graphs (Zhang et al., 2023b; Fang et al., 2023a). Formally, let $P_{\mathcal{G}}$ denote the distribution of training graphs, and $P_{{\mathcal{G}}^{\prime}}$ represents the distribution of explanation subgraphs. The core issue in the GIB objective function is caused by $P_{\mathcal{G}}\neq P_{{\mathcal{G}}^{\prime}}$. This distributional disparity undermines the predictive reliability of the model $f(\cdot)$, trained on $P_{\mathcal{G}}$ when applied to subgraphs from $P_{{\mathcal{G}}^{\prime}}$. As a result, the predictive power of explanations provided by $f(G^{\prime})$ is unreliable approximation of $P(Y|G^{\prime})$ in  equation 3.
@@ 3 Graph Information Bottleneck with Proxy Graphs
In this section, we propose the concept of a proxy graph to mitigate the above-mentioned OOD issue. This proxy graph not only retains the label information present in $G^{\prime}$ but also conforms to the distribution of the original graph dataset. Specifically, we assume that proxy graphs $\tilde{G}$ are drawn from a distribution $P_{\mathcal{G}}$ and reformulate the estimation of $P(Y|G^{\prime})$ by marginalizing over the distribution of proxy graphs as follows.
$P(Y|G^{\prime})=\mathbb{E}_{\tilde{G}\sim P_{\mathcal{G}}}[P(Y|\tilde{G})\cdot P(\tilde{G}|G^{\prime})]$ (4)
@@
Given the combinatorial complexity inherent in graph structures, it is computationally infeasible to enumerate all potential proxy graphs $\tilde{G}$ from the unknown distribution $P_{\mathcal{G}}$, which is necessary for calculating equation 5. To overcome this challenge, we propose approximating $P_{\mathcal{G}}$ with the parameterized function $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$. Consequently, we estimate equation 5 by sampling proxy graphs from $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$, leading to the following formulation:
@@
The first constraint ensures that $\tilde{G}$ is sampled from a distribution that approximates $P_{\mathcal{G}}$, effectively addressing the OOD challenge. The second constraint guarantees that the label information preserved in $\tilde{G}$ is similar to that in $G^{\prime}$. Therefore, our proxy graph-induced objective function becomes:
@@
Bi-level optimization In equation 8, we formulate a joint optimization loss function that aims to identify the optimal explanation alongside its corresponding proxy graphs. Building upon this, we refine the framework by replacing the first constraint with a distributional distance measure, specifically the Kullback-Leibler (KL) divergence, between $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$ and $P_{\mathcal{G}}$. This leads to the development of a bi-level optimization model. Formally, the model is expressed as follows:
$\operatorname*{arg\,min}_{G^{\prime}}{I(G,G^{\prime})-\alpha\mathbb{E}_{G^{\prime},Y}[\log\mathbb{E}_{\tilde{G}\sim Q_{\bm{\phi}^{*}}(\tilde{G}|G^{\prime})}[P(Y|\tilde{G}^{*})]]}$ (9)
@@ 3.1 Derivation of Outer Optimization
For the outer optimization objective, we elaborate on operationalizing $I(G,G^{\prime})$. Akin to the original graph, We denote the explanation subgraph $G^{\prime}$ by $({\mathcal{V}},{\mathcal{E}}^{\prime},{\bm{X}})$, whose adjacency matrix is ${\bm{A}}^{\prime}$. We follow (Miao et al., 2022) to include a variational approximation distribution $R(G^{\prime})$ for the distribution $P(G^{\prime})$. Then, we obtain its upper bounds as follows.
$I(G,G^{\prime})\leq\mathbb{E}_{G}[\text{KL}(P(G^{\prime}|G)||R(G^{\prime}))]$ (10)
We follow the Erdős–Rényi model (Erdős et al., 1960) and assume that each edge in $G^{\prime}$ has a probability of being present or absent, independently of the other edges. Specifically, we assume that the existence of an edge $(u,v)$ in $G^{\prime}$ is determined by a Bernoulli variable $A_{u,v}^{\prime}\sim\text{Bern}(\pi_{uv})$. Thus, $P(G^{\prime}|G)$ can be decompose with $P(G^{\prime}|G)=\prod_{(u,v)\in{\mathcal{E}}}P(A_{uv}^{\prime}|\pi_{uv})$. The bound is always true for any prior distribution $R(G^{\prime})$. We follow an existing work and assume that in prior distribution, the existence of an edge $(u,v)$ in $G^{\prime}$ is determined by another Bernoulli variable $A^{{}^{\prime\prime}}_{u,v}\sim\text{Bern}(r)$, where $r\in[0,1]$ is a hyper-parameter (Miao et al., 2022), independently to the graph $G$. We have $R(G^{\prime})=P(|{\mathcal{E}}|)\prod_{(u,v)\in{\mathcal{E}}}P(A^{{}^{\prime\prime}}_{uv})$. Thus, the KL divergence between $P(G^{\prime}|G)$ and the above marginal distribution $R(G^{\prime})$ becomes:
@@
$=$ $\sum_{(u,v)\in{\mathcal{E}}}\pi_{uv}\log\frac{\pi_{uv}}{r}+(1-\pi_{uv})\log\frac{1-\pi_{uv}}{1-r}+\text{Const}.$
By replacing $I(G^{\prime}G^{\prime})$ with the above tractable upper bound, we obtain the loss function, denoted by $\mathcal{L}_{\text{exp}}$, to train the explainer $\Psi(\cdot)$.

### v2

# Generating In-Distribution Proxy Graphs for Explaining Graph Neural Networks

@@ Abstract
Graph Neural Networks (GNNs) have become a building block in graph data processing, with wide applications in critical domains. The growing needs to deploy GNNs in high-stakes applications necessitate explainability for users in the decision-making processes. A popular paradigm for the explainability of GNNs is to identify explainable subgraphs by comparing their labels with the ones of original graphs. This task is challenging due to the substantial distributional shift from the original graphs in the training set to the set of explainable subgraphs, which prevents accurate prediction of labels with the subgraphs. To address it, in this paper, we propose a novel method that generates proxy graphs for explainable subgraphs that are in the distribution of training data. We introduce a parametric method that employs graph generators to produce proxy graphs. A new training objective based on information theory is designed to ensure that proxy graphs not only adhere to the distribution of training data but also preserve explanatory factors. Such generated proxy graphs can be reliably used to approximate the predictions of the labels of explainable subgraphs. Empirical evaluations across various datasets demonstrate our method achieves more accurate explanations for GNNs.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a pivotal technology for handling graph-structured data, demonstrating remarkable performance in various applications including node classification and link prediction (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Scarselli et al., 2008). Their growing use in critical sectors such as healthcare and fraud detection has escalated the need for explainability in their decision-making processes (Wu et al., 2022; Li et al., 2022; Zhang et al., 2024). To meet this demand, a variety of explanation methods have been recently developed to interpret the behavior of GNN models. These methods primarily concentrate on identifying a subgraph that significantly impacts the model’s prediction for a particular instance (Ying et al., 2019; Luo et al., 2020).
@@
Figure 1: Examples of the explanation process and out-of-distribution problem. (a) is the explanation process for a graph learning model. The original graph $G$ undergoes an explanation process, resulting in an explanation graph $G^{\prime}$ that highlights the most significant features and relationships; (b) shows the explanation graph $G^{\prime}$ is out of distribution where the GNN is trained.
Although it is intuitively correct, the underlying assumption of the aforementioned approach – GNN model can make accurate predictions on explanation subgraphs – may not always hold. As shown in Figure 1, explanation subgraphs can significantly deviate from the distribution of original graphs, leading to an Out-Of-Distribution (OOD) issue (Zhang et al., 2023b; Fang et al., 2023c, a; Zheng et al., 2024). For instance, in the MUTAG dataset (Debnath et al., 1991), each graph represents a molecule, with nodes symbolizing atoms and edges indicating chemical bonds. The molecular graphs in this dataset usually contain hundreds of edges. In contrast, the $NO_{2}$ functional group, identified as a key subgraph influencing positive mutagenicity in a molecule, comprises merely 2 edges. This stark contrast in structural properties leads to a significant difference in the distributions of explanation subgraphs and original graphs. Since the model was trained with original graphs, the reliability of predictions on subgraphs is undermined due to the distribution shifting problem.
Several pioneering studies have attempted to address this distributional challenge (Fang et al., 2023c; Zhang et al., 2023b, a). For example, CGE regards the GNN model as a transparent, fully accessible system. It considers the GNN model as a teacher network and employs an additional “student” network to predict the labels of explanation subgraphs (Fang et al., 2023c). As another example, MixupExplainer (Zhang et al., 2023b) generates a mixed graph for the explanation by blending it with a non-explanatory subgraph from a different input graph. This method posits that the mixup graph aligns with the distribution of the original input graphs. However, this claim is predicated on a rather simplistic assumption that the explanation and non-explanatory subgraphs are independently drawn. However, in real-world applications, these methods often face practical limitations. The dependence on a “white box” model in CGE and the oversimplified assumptions in MixupExplainer are not universally applicable. Instead, GNN models are usually given as “black boxes”, and the graphs in real-life applications do not conform to strict independence constraints, highlighting the need for more versatile and realistic approaches to the OOD problem.
@@
  • We systematically analyze and address the challenge of the OOD issue in explainable GNNs, which is pivotal for enhancing the reliability and interpretability of GNNs in real-world applications.
@@
  • Through comprehensive experiments on various real-world datasets, we substantiate the effectiveness of our proposed approach, showcasing its practical utility and superiority in producing explanations.
@@ 2 Notations and Preliminary
2.1 Notations and Problem Formulation
We denote a graph $G$ from an graph set ${\mathcal{G}}$ by a triplet $({\mathcal{V}},{\mathcal{E}};{\bm{X}})$, where ${\mathcal{V}}=\{v_{1},v_{2},...,v_{n}\}$ is the node set and ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the edge set. ${\bm{X}}\in{\mathbb{R}}^{n\times d}$ is the node feature matrix, where $d$ is the feature dimension and the $i$-th row is the feature vector associated with node $v_{i}$. The adjacency matrix of $G$ is denoted by ${\bm{A}}\in\{0,1\}^{n\times n}$, which is determined by the edge set $\mathcal{E}$ such that ${A}_{ij}=1$ if $(v_{i},v_{j})\in\mathcal{E}$, ${A}_{ij}=0$, otherwise. In this paper, we focus on the graph classification task, and node classification can be converted to computation graph classification problem (Ying et al., 2019; Luo et al., 2020). Specifically, for graph classification task, each graph $G$ is associated with a label $Y\in\mathcal{Y}$. The to-be-explained GNN model $f(\cdot)$ has been well-trained to classify $G$ into its class, i.e., $f:{\mathcal{G}}\mapsto\{1,2,\cdots,|\mathcal{Y}|\}$.
Following the existing works (Ying et al., 2019; Luo et al., 2020; Yuan et al., 2022; Huang et al., 2024), the explanation methods under consideration in this paper are model/task agnostic and treat GNN models as black boxes — i.e., the so-called post-hoc, instance-level explanation methods. Formally, our research problem is described as follows:
@@ 2.1 Notations and Problem Formulation
Given a to-be-explained GNN model $f(\cdot)$ and a set of graphs ${\mathcal{G}}$, the goal of post-hoc instance-level explanation is to learn a parametric function so that for an arbitrary graph $G\in{\mathcal{G}}$, it finds a compact subgraph $G^{*}\subseteq G$ that can “explain” the prediction $f(G)$. The parametric mapping $\Psi_{\psi}:{\mathcal{G}}\mapsto{\mathcal{G}}^{*}$ is called an explanation function, where ${\mathcal{G}}^{*}$ is the alphabet of $G^{*}$, and $\psi$ is the parameter of the explanation function.
2.2 Graph Information Bottleneck as the Objective Function
The Information Bottleneck (IB) principle, foundational in learning dense representations, suggests that optimal representations should balance minimal and sufficient information for predictions (Tishby et al., 2000). This concept has been adapted for GNNs through the Graph Information Bottleneck (GIB) approach, consolidating various post-hoc GNN explanation methods like GNNExplainer (Ying et al., 2019) and PGExplainer (Luo et al., 2020). The GIB framework aims to find a subgraph $G^{*}$ to explain a GNN model’s prediction on a graph $G$ as (Zhang et al., 2023b):
@@ 2.2 Graph Information Bottleneck as the Objective Function
where $G^{\prime}$ is a candidate explanatory subgraph, $Y$ is the label, and $\alpha$ is a balance parameter. The first term is the mutual information between the original graph $G$ and the explanatory subgraph $G^{\prime}$ and the second term is negative mutual information of $Y$ and the explanatory subgraph $G^{\prime}$. At a high level, the first term encourages detecting a small and dense subgraph for explanation, and the second term requires that the explanatory subgraph is label preserving.
Due to the intractability of mutual information between $Y$ and $G^{\prime}$ in equation 1, some existing works (Wu et al., 2020; Miao et al., 2022) derived a parameterized variational lower bound of $I(Y,G^{\prime})$:
@@
The OOD Problem in GIB. In the existing research, the estimation of $P(Y|G^{\prime})$ is typically achieved by applying the to-be-explained model $f(\cdot)$ to the input graph $G^{\prime}$ (Ying et al., 2019; Miao et al., 2022). A critical assumption in these approaches involves the GNN model’s ability to accurately predict candidate explanation subgraphs $G^{\prime}$. This assumption, however, overlooks the OOD problem, where the distribution of explanation subgraphs significantly deviates from that of the original training graphs (Zhang et al., 2023b; Fang et al., 2023a, c, b; Zheng et al., 2024; Amara et al., 2023). Formally, let $P_{\mathcal{G}}$ be the distribution of the training graphs, and $P_{{\mathcal{G}}^{\prime}}$ be the distribution of the explanation subgraphs. The core issue in the GIB objective function is caused by $P_{\mathcal{G}}\neq P_{{\mathcal{G}}^{\prime}}$. This distributional disparity undermines the predictive reliability of the model $f(\cdot)$, trained on $P_{\mathcal{G}}$ when applied to subgraphs from $P_{{\mathcal{G}}^{\prime}}$. As a result, the predictive power of explanations provided by $f(G^{\prime})$ is an unreliable approximation of $P(Y|G^{\prime})$ in  equation 3.
@@ 3 Graph Information Bottleneck with Proxy Graphs
In this section, we propose the concept of proxy graphs to mitigate the aforementioned OOD issue. A proxy graph not only retains the label information present in $G^{\prime}$ but also conforms to the distribution of the original graph dataset. Specifically, we assume that proxy graphs $\tilde{G}$ are drawn from the distribution $P_{\mathcal{G}}$ and reformulate the estimation of $P(Y|G^{\prime})$ by marginalizing over the distribution of proxy graphs as follows:
$P(Y|G^{\prime})=\mathbb{E}_{\tilde{G}\sim P_{\mathcal{G}}}[P(Y|\tilde{G})\cdot P(\tilde{G}|G^{\prime})].$ (4)
@@
Given the combinatorial complexity inherent in graph structures, it is computationally infeasible to enumerate all potential proxy graphs $\tilde{G}$ from the unknown distribution $P_{\mathcal{G}}$, which is necessary for calculating equation 5. To overcome this challenge, we propose to approximate $P_{\mathcal{G}}$ with the parameterized function $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$, and estimate equation 5 with a Monte Carlo estimator that samples proxy graphs from $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$. That is
@@
where the first constraint ensures that $\tilde{G}$ is sampled from a distribution that approximates $P_{\mathcal{G}}$, effectively addressing the OOD challenge. The second constraint guarantees that the label information preserved in $\tilde{G}$ is similar to that in $G^{\prime}$. Therefore, combining equation 3, equation 6 and equation 7, our proxy graph-induced objective function becomes:
@@
Bi-level optimization. In equation 8, we formulate a joint optimization loss function that aims to identify the optimal explanation alongside its corresponding proxy graphs. Building upon this, we refine the framework by replacing the first constraint with a distributional distance measure, specifically the Kullback-Leibler (KL) divergence, between $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$ and $P_{\mathcal{G}}$. This leads to the development of a bi-level optimization model. Formally, the model is expressed as follows:
$\operatorname*{arg\,min}_{G^{\prime}}{I(G,G^{\prime})-\alpha\mathbb{E}_{G^{\prime},Y}[\log\mathbb{E}_{\tilde{G}\sim Q_{\bm{\phi}^{*}}(\tilde{G}|G^{\prime})}[P(Y|\tilde{G})]]}$ (9)
@@ 3.1 Derivation of Outer Optimization
For the outer optimization objective, we elaborate on instantiating the first term $I(G,G^{\prime})$. Akin to the original graph, we denote the explanation subgraph $G^{\prime}$ by $({\mathcal{V}},{\mathcal{E}}^{\prime},{\bm{X}})$, whose adjacency matrix is ${\bm{A}}^{\prime}$. We follow (Miao et al., 2022) to include a variational approximation distribution $R(G^{\prime})$ for the distribution $P(G^{\prime})$. Then, we obtain its upper bound as follows:
$I(G,G^{\prime})\leq\mathbb{E}_{G}[\text{KL}(P(G^{\prime}|G)||R(G^{\prime}))].$ (10)
We follow the Erdős–Rényi model (Erdős et al., 1960) and assume that each edge in $G^{\prime}$ has a probability of being present or absent, independently of the other edges. Specifically, we assume that the existence of an edge $(u,v)$ in $G^{\prime}$ is determined by a Bernoulli variable $A_{u,v}^{\prime}\sim\text{Bern}(\pi^{\prime}_{uv})$. Thus, $P(G^{\prime}|G)$ can be decomposed as $P(G^{\prime}|G)=\prod_{(u,v)\in{\mathcal{E}}}P(A_{uv}^{\prime}|\pi^{\prime}_{uv})$. The bound is always true for any prior distribution $R(G^{\prime})$. We follow an existing work and assume that in the prior distribution, the existence of an edge $(u,v)$ in $G^{\prime}$ is determined by another Bernoulli variable $A^{{}^{\prime\prime}}_{u,v}\sim\text{Bern}(r)$, where $r\in[0,1]$ is a hyper-parameter (Miao et al., 2022), independently to the graph $G$. We have $R(G^{\prime})=P(|{\mathcal{E}}|)\prod_{(u,v)\in{\mathcal{E}}}P(A^{{}^{\prime\prime}}_{uv})$. Thus, the KL divergence between $P(G^{\prime}|G)$ and the above marginal distribution $R(G^{\prime})$ becomes:
@@
$=$ $\sum_{(u,v)\in{\mathcal{E}}}\pi^{\prime}_{uv}\log\frac{\pi^{\prime}_{uv}}{r}+(1-\pi^{\prime}_{uv})\log\frac{1-\pi^{\prime}_{uv}}{1-r}+\text{Const}.$
By replacing $I(G,G^{\prime})$ with the above tractable upper bound, we obtain the loss function, denoted by $\mathcal{L}_{\text{exp}}$, for training the explainer $\Psi(\cdot)$.